Ms. Teresa Monson is employed by Elmwood Inc. She has asked the employer for a $300,000 interest free loan that will be used to acquire a summer cottage in Huntsville, Ontario. The cottage will be used exclusively as a recreational property. As she is a highly valued employee, Elmwood Inc. is considering her request. Ms. Monson can acquire a regular mortgage at a rate of 4.5 percent. Assume that the relevant prescribed rate is 2 percent for all periods that the employee loan will be outstanding . Ms. Monson's tax rate on any additional income is 46 percent . Elmwood Inc. has alternative investment opportunities that earn a before tax rate of 7 percent . Elmwood Inc. is subject to a tax rate of 28 percent on additional amounts of income . Required : Evaluate Ms. Monson's suggestion of providing her with an interest free loan in lieu of salary from the point of view of the cost to Elmwood Inc.
Answer #1

alternative 1

total interest rate =4.5+2=6.5

principal amount=300000

interest amount =300000*6.5%=19500

interest after tax=19500(1-.46)=8969

alternative 2

interest rate=7%

interest amount=300000*7%=21000

interest after tax=21000(1-.28)=5879

so the interest after tax is more in second alternative.